OpenStudy (anonymous):

Can someone please help me find the derivative of this problem? I will fan and medal.

OpenStudy (jhannybean):

Whats the function?

OpenStudy (anonymous):

I'm typing it

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }\sin^\frac{ 3 }{ 2 }x-\frac{ 2 }{ 7 }\sin^\frac{ 7 }{ 2 }x\]

OpenStudy (jhannybean):

\[\frac{2}{3}\left(\sin^{2/3}x\right)' \iff \frac{2}{3}\left((\sin(x))^{2/3}\right)'\]

OpenStudy (wmj259):

If you look carefully you see that the two functions are connected with the subtract sign. So you can break that function into 2 separate derivatives.

OpenStudy (wmj259):

\[\frac{ d }{ dx } (\frac{ 2 }{ 3 }\sin ^{2/3}(x)) - \frac{ d }{ dx }((\frac{ 2 }{ 7 }\sin ^{7/2}x))\]

OpenStudy (jhannybean):

Use the power rule, then the chain rule.

OpenStudy (wmj259):